OpenStudy (anonymous):

George measured the capacity of a cardboard box as 400 in3. The true capacity of the cardboard box is 450 in3. What is the percent error in George’s measurement? If necessary, round to the nearest percent.

jimthompson5910 (jim_thompson5910):

what is the difference in the two values? ie how far apart are they?

OpenStudy (anonymous):

50 right?

jimthompson5910 (jim_thompson5910):

yes, 450 - 400 = 50

jimthompson5910 (jim_thompson5910):

divide that difference by the true measurement 50/450 = ??

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

roughly 0.1111, yes

jimthompson5910 (jim_thompson5910):

multiply 0.1111 by 100 to turn it into a percentage 0.1111*100 = 11.11%

jimthompson5910 (jim_thompson5910):

so George is off by roughly 11.11%

jimthompson5910 (jim_thompson5910):

oh " If necessary, round to the nearest percent." so 11.11% turns into 11%

OpenStudy (anonymous):

thank you if i have anymore question i will tag you :)